Grade 11Physical Chemistry

If 80g of X combines with 1.5 X 10^23 atoms of Y to form X2 without any of either element remaining. Determine gram atomic weight of X

ekta vishwakarma
9 Years agoGrade 11

1 Answer

Rohit
9 Years ago
X2Y expresses molar ratio i.e Y=1.5 × 10 ^23 atomsSo X = 2(1.5 × 10^23) Therefore When mass of X is 80 gm . No. Of atoms is =3 × 10^23 atomsGram atomic weight is the weight 1 mole of X and 1 mole = 6 × 10^23 atoms By cross multiplication u will get atomic at as 160 grams